Night report 2021-04-09

Observers: Priscila Pessi

Support: Christian Vogl, Andreas Floers, Geza Csoernyei, Alexander Holas and Stefan Taubenberger

TAT: Giuliano Pignata, Jacob Teffs, Michel Dennefeld

EFOSC Calibrations:

Bias, dome_flats, arcs.

Skyflats UBVR (last R flats are not usefull because the stars appeared)

Weather:

Start: humidity 51%, wind 4.9 m/s, seeing 1.3

Middle: humidity 18% wind 7.3 m/s, seeing 1.05

End: humidity 21% wind 6.4 m/s, seeing 1.22

Comments:

23:13 UT image analysis.

05:00 UT, CCD had a problem the OB was aborted although the integration was finished.

EFOSC

SOFI: Switch at 05:58 UT, image analysis and focus after switch.

One of the quadrants was a bit noisy, we did our best to reduce the noise to the maximum possible.